Memoe Plc is a large firm with a market capitalisation of €3,560,000,000 (3.56 Billion). The firm has one billion ordinary shares issued.

The firm’s long-term debt consists of a single bond issue of 100 million bonds with a current market price of €92.60 per bond. These bonds are redeemable at par of €100 per bond in 5 years’ time. The coupon on these bonds is 6.86% paid annually. Corporate tax rate is 12.5%.

You have calculated the firm’s beta using historical data over a five-year period as 1.36. the risk-free rate of interest is 4.25% per annum, the returns on equity markets over the past five years average 7.3%.

a) Calculate the Weighted Average Cost of Capital
b) Comment on your result. Do you think that the Weighted Average Cost of Capital is realistic? How risky is this firm compared to the average?
c) What is the company’s current share price?
Answer #1

Weight of equity = (3.56/3.56+9.26)=0.27

Weight of debt = (9.26/12.82)=0.73

Cost of equity = risk-free rate + (Beta x market risk premium)

4.25+(1.36*7.3)= 14.18%

WACC = [weight of debt x cost of debt x (1 - tax rate)] + (weight of equity x cost of equity)

=(0.73*6.86*0.875)+(0.27*14.18)

=4.38+3.83

= 8.21%

WACC represents the expense of raising one additional dollar of money. A WACC of 8.21% means the company must pay its investors an average of $0.0821 in return for every $1 in extra funding. It can definitely be realistic depending on other factors. But as weight of debt is significantly high, company is definitely at a considerable risk.

Share Price = 3.56 Billion/ 1 Billion = \epsilon 3.56.